# Simulating many-engine spacecraft: Exceeding 1 quadrillion degrees of freedom via information geometric regularization

### v3

@@
Abstract
We present an optimized implementation of the recently proposed information geometric regularization (IGR) for unprecedented scale simulation of compressible fluid flows applied to multi-engine spacecraft boosters. We improve upon state-of-the-art computational fluid dynamics (CFD) techniques along computational cost, memory footprint, and energy-to-solution metrics. Unified memory on coupled CPU–GPU or APU platforms increases problem size with negligible overhead. Mixed half/single-precision storage and computation on well-conditioned numerics is used. We simulate flow at 200 trillion grid points and 1 quadrillion degrees of freedom, exceeding the current record by a factor of 20. A factor of 4 wall-time speedup is achieved over optimized baselines. Ideal weak scaling is seen on OLCF Frontier, LLNL El Capitan, and CSCS Alps using the full systems. Strong scaling is near ideal at extreme conditions, including 80% efficiency on CSCS Alps with an 8 node baseline and stretching to the full system.
Index Terms:
@@ Abstract
I Justification for ACM Gordon Bell Prize
Largest compressible computational fluid dynamics simulation, exceeding 1 quadrillion degrees of freedom and a factor of 20 beyond state-of-the-art. Enabled by a large-scale optimized use of inviscid (information geometric) regularization. Time- and energy-to-solution decrease by up to a factor of 4 and 5.4 in double precision, more in single and mixed precision.
II Performance Attributes
TABLE I: Summary of Performance Attributes
@@ II Performance Attributes
III Overview of the Problem
The 21st century witnesses a “new space race” [1] driven by private companies replacing government agencies in providing launch services. The resulting drop in launch cost enabled many business models, ranging from satellites to space manufacturing [2]. The new emphasis on cost efficiency motivated rocket design innovations like reusing rocket stages and leveraging economies of scale.
@@ III Overview of the Problem
Large arrays of small engines create new design challenges. Interacting exhaust plumes of densely packed engines can propel hot gas toward the rocket base, heating it. This so-called base heating can cause mission failure [3], mandating heat shields on the rocket base that increase weight and cost.
Mitigating base-heating most cost-effectively, in terms of both dollars and weight, requires appreciating the mechanism by which engine exhaust is reflected towards the rocket and which parts are most affected. A key difficulty in experimental approaches to understanding plume recirculation is that it depends on numerous parameters, including varying ambient pressure as the rocket traverses the atmosphere and engine thrust vectoring for steering. A detailed flow field characterization under a broad range of conditions is only feasible with numerical simulations. They even allow probing the impacts of changes to the rocket design. Prior work on the numerical simulation of interacting rocket plumes was limited to small numbers (up to 7) of rocket engines and limited resolution (up to $\approx{}10$ million grid points) [3, 4, 5]. Our work addresses this shortcoming by utilizing the latest flagship exascale systems, leveraging their hardware design and coupling novel algorithms, computational methods, and the optimizations they enable. With this, we can simulate the interaction of rocket engine plumes at unprecedented scales.
@@
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})$ $=0,$ (1)
$\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u}+p\mathbf{I}-\boldsymbol{\tau})$ $=\mathbf{0},$ (2)
$\frac{\partial E}{\partial t}+\nabla\cdot\left[(E+p)\mathbf{u}-\mathbf{u}\cdot\boldsymbol{\tau}\right]$ $=0,$ (3)
@@
$p=(\gamma-1)\rho e,\quad\text{for}\quad e\coloneqq E/\rho-\|\mathbf{u}\|^{2}/2$ (4)
@@
$\tau_{ij}=\mu\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right)+\left(\zeta-{2}\mu/{3}\right)\delta_{ij}\frac{\partial u_{i}}{\partial x_{j}}.$ (5)
Equations˜1, 2 and 3 describe the conservation of mass, momentum, and energy, and $\boldsymbol{\tau}$ is the viscous stress tensor and $p$ the pressure. Tracking mixture ratios of different gases/fluids and their chemical reactions would be a natural extension of the demonstration. For this work and its implications, we focus on eqs.˜1, 2, 3, 4 and 5.
@@
Figure˜1 shows a visualization for a simulation run on CSCS Alps. The visualization shows the interactions between an array of 33 Mach 10 rocket engines organized in an array inspired by that of the SpaceX superheavy booster, each resolved by 600 grid cells across its outlet. The high resolution allows fine-scale details to be represented despite the high Mach number. The simulation of fig.˜1 uses a rectilinear grid of 3.3T cells and ran for 16 hours on 9.2K GH200s (2300 Alps nodes).
Figure 1: Simulation results showing the interacting plumes from an array of 33 thrusters in a configuration inspired by the SpaceX superheavy booster with 16.5 trillion degrees of freedom. (The thrusters themselves are for visualization purposes. We model them through inflow boundary conditions.)
IV Current State of the Art
IV-A Shock capturing
@@
Shock waves are the dominant concern of high-speed compressible fluid dynamics simulation. They arise in numerous natural and man-made phenomena, including supernovae, air, and spacecraft. The velocity and density fields of compressible high-speed flows sharpen over time, eventually forming macroscopic discontinuities or shock waves. However, on the microscopic scale, the gas viscosity balances the steepening of shock waves, resulting in smooth profiles. In practice, this happens on the scale of the mean free path of gas particles, orders of magnitude smaller than the quantities of interest.
@@
Higher-order numerical methods exploit the target solution’s regularity to approximate it with smooth functions, most frequently polynomials. The mean-free path is many orders of magnitude shorter than a realistic computational grid spacing. Thus, shocks appear as discontinuities on the grid scale, and the direct application of higher-order methods leads to Gibbs–Runge oscillations and subsequent simulation failure. Shock capturing modifies either the equation or its discretization to obtain a well-behaved object on the grid scale. It amounts to coarse-graining the microscopic shock and correctly representing its macroscopic effects without grid-resolving it. Crucially, the resulting coarse-scale model should preserve smooth grid-scale oscillations due to turbulence, reactions, or acoustic effects.
@@
Artificial viscosity mitigates Gibbs–Runge oscillations at the cost of excessive dissipation of fine-scale features. To remedy the latter, numerous approaches apply artificial diffusivity adaptively, in the vicinity of the shock [6, 7, 8].
@@
Realizing a viscous regularization is challenging in practice. A lack of viscosity creates spurious oscillations, while excessive viscosity dissipates the solution. Common choices, such as the localized artificial diffusivity (LAD) of [6], attempt to strike this balance. Viscous methods spread the shock over multiple grid points, but the resulting shock profile is not high-order smooth (fig.˜2 (a,i)). This can lead still to Gibbs–Runge oscillations overcoming the regularization and destabilizing solutions. Increasing the shock width requires increasing the strength of the artificial viscosity, which smears true physical features critical to representing the flow (fig.˜2, (b,i)). In the presence of sufficiently strong shocks, the required artificial viscosity affects the CFL numbers of the explicit time steppers considered state-of-the-art for such hyperbolic problems.
@@
Limiters are an alternative to artificial viscosity. They adaptively lower the order of the numerical method at shocks [9, 10]. They are more robust but also risk dissipating fine-scale features. Riemann solvers aim to mitigate this problem but add computational cost [11, 12, 13].
IV-B GPU memory
@@
The evolution of GPU memory over the past decades, in terms of capacity and bandwidth, has been highly beneficial for CFD applications, which exhibit low arithmetic intensity. Thus, performance is limited by memory bandwidth. The spatio-temporal scale separation in fluid flows requires a high resolution, necessitating large amounts of memory to be accessed at high speeds. Both NVIDIA and AMD have blurred the lines between GPU and CPU memory by introducing coherent CPU-to-GPU interfaces: $900\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ bidirectional bandwidth for the NVIDIA Grace Hopper and InfinityFabric at $8\times 72\,$\mathrm{GB}\text{/}\mathrm{s}$\$ for the AMD Trento+4 MI250X configuration of Frontier. These fast interconnects, along with technologies such as unified virtual memory (UVM), allow CPU memory for GPU computations and even (in the case of Grace Hopper) allow the GPU to saturate the host memory bandwidth. Further still, AMD introduced the MI300A in LLNL’s El Capitan and Tuolumne, with a single physical HBM pool accessed by both CPU and GPU devices. This strategy avoids concern about device and host pointers or memory store location.
IV-C Floating point computation
@@
The advent of artificial intelligence fuels the development of algorithms with reduced and mixed precision, though their adoption in HPC applications remains limited. Most traditional scientific applications rely on FP64 computation and storage. However, recent advances in software and numerical methods demonstrate the viability of sub-FP64 precision for well-conditioned numerical methods and a range of traditional fluid flow problems [14]. Still, FP64 is the de facto standard for compressible, shock-laden flow simulation and traditional shock-capturing techniques. WENO-based reconstruction methods and approximate Riemann solves, considered state-of-the-art and used as the baseline in this work, involve poorly conditioned operations and thus are not well suited to reduced precision [15]. We elide such issues herein by avoiding numerical shock capturing entirely via IGR.
IV-D Large CFD simulations
@@
Previous Gordon Bell winner [16] performed a 10T grid point CFD simulation on IBM BlueGene/Q. The system’s substantial CPU memory enabled this simulation. However, the wall time required to process such a large simulation makes achievable physical time scales short. More recently, [17] solved a compressible CFD problem of 10T grid points when extrapolating to 100% of Frontier.
@@
The current largest accelerator-based flow simulations are for incompressible flows with fewer degrees of freedom. The largest simulation represented turbulence on a 30T point grid using OLCF Frontier [18]. Large time step costs arise from all-to-all communications that dominate time-to-solution.
V Innovations Realized
V-A The need for scale
@@
Figure 2: Inviscid regularization. Localized artificial diffusion (LAD) spreads shocks over a user-defined width (a,i). The resulting curve is not high-order smooth. This can cause methods with a high order of accuracy to develop oscillations and, ultimately, fail. Increasing the width for coarser discretizations yields unphysical and significant dissipation of oscillatory solution profiles (b,i). Information geometric regularization replaces shocks with smooth profiles (a,ii) at the grid scale and preserves oscillatory features (b,ii).
V-B Shock treatment with information geometric regularization
We avoid the limitations of conventional shock-capturing approaches via the first inviscid regularization of the equations, called information geometric regularization (IGR) as recently proposed by Cao and Schäfer [19]. IGR was first derived in the pressureless (infinite Mach number) case, where shocks amount to the loss of injectivity of the flow map $x\mapsto\phi_{t}(x)$ that maps gas particles from their initial position to their position at time $t$. IGR modifies the geometry according to which the flow map evolves, such that particle trajectories $t\mapsto\phi_{t}(x_{0})$ do not cross and instead asymptotically approach each other (see fig.˜3). This preserves the long-time post-shock behavior, prevents the formation of grid-level singularities, and recovers the nominal vanishing viscosity solutions in the limit. This result has been proven rigorously in the unidimensional, pressureless case [20].
@@ V-B Shock treatment with information geometric regularization
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})$ $=0,$ (6)
$\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u}+(p+\Sigma)\mathbf{I}-\boldsymbol{\tau})$ $=\mathbf{0},$ (7)
$\frac{\partial E}{\partial t}+\nabla\cdot\left[(E+p+\Sigma)\mathbf{u}-\mathbf{u}\cdot\boldsymbol{\tau}\right]$ $=0,$ (8)
$\alpha\left[\operatorname{tr}{(\nabla\mathbf{u})^{2}}+\operatorname{tr}^{2}{(\nabla\mathbf{u})}\right]$ $=\frac{\Sigma}{\rho}-\alpha\nabla\!\cdot\!\left(\frac{\nabla\Sigma}{\rho}\right)\!.$ (9)
@@
Figure 3: Information geometric regularization modifies shocks by changing the geometry according to which the flow map $\phi_{t}$ evolves in time. In the modified geometry, the trajectories of two tracer particles $t\mapsto\phi_{t}(x_{1}),\phi_{t}(x_{2})$ converge in $t$ rather than cross. The regularization strength $\alpha$ determines the rate of convergence. The vanishing viscosity solution is recovered in the $\alpha\rightarrow 0$ limit. Figure adapted from Cao and Schäfer [19] with author permission.
@@
IGR allows us to bypass shock capturing and directly use a third or fifth-order accurate finite volume method with Lax–Friedrichs numerical fluxes for the hyperbolic part of the equation. Due to the high Reynolds number of the problem under consideration, we find that a second-order accurate approximation of the derivatives of $\mathbf{u}$ suffices to compute the viscous stress tensor. We reuse these derivatives to compute the left side of eq.˜9 and discretize the elliptic operator on the right using a standard 7-point stencil.
@@
For each computation of the hyperbolic flux, we solve eq.˜9 using up to 5 sweeps of Jacobi or Gauss–Siedel iteration, with the previously computed $\Sigma$ as an initial guess. We use a third-order accurate Runge–Kutta time stepper [21]. For a single species (advected fluid) case, the total number of floating point numbers stored by our scheme is $17N+o(N)$, where $o(N)$ is the number of grid points. This includes $5$ state variable arrays (density, energy, and three momenta), which are the degrees of freedom of the solution per grid cell. We also hold $5$ arrays for the Runge–Kutta sub-step, $5$ arrays for the right-hand side of the discretized PDE system, one array for $\Sigma$, and one for the right-hand side of eq.˜9. An additional copy of $\Sigma$ is required if Jacobi sweeps are used for the iterative solve.
V-C The algorithm
V-D Optimizations
Our implementation eliminates the storage of the reconstructed states (algorithms 1 to 1), velocity gradients (algorithms 1 to 1), and fluxes (algorithms˜1, 1 and 1) and keeps all operations in a single kernel (algorithm˜1). The memory footprint is reduced by storing the intermediate variables as thread-local temporary arrays within this kernel. The algorithm only requires storing $2$ copies of the conservative variables, the net flux at each grid point, and the solution and left side of the elliptic PDE in eq.˜9
Each thread solves an approximate Riemann problem at the grid cell–cell interface and accumulates its contribution to the right-hand side at overlapping cells via atomic operations to prevent race conditions. During the reconstruction along the 1st coordinate dimension ($x$, here), the contributions to the left hand side of equation eq.˜9 are computed using the velocity gradients for the viscous fluxes (algorithm˜1). The elliptic PDE in eq.˜9 is solved after its left side is computed (algorithm˜1). The entropic pressure added directly to the flux in the 2nd and 3rd dimensions (algorithms 1 to 1). The flux contribution of the entropic pressure in the first dimension is completed separately after the elliptic solve (algorithms 1 to 1). The above decreases memory use $25$-fold compared to an optimized 5th-order accurate WENO and Riemann solver implementation in the same codebase [22].
TABLE II: Node and full system properties of the supercomputers tested in this work. TOP500 rankings from June 2025.
Figure 4: Schematic of the chip-to-chip (C2C) transfers of intermediate time-step variables between the on-node CPU and GPU devices. The time sub-steps are $q_{1,2}$ and the full step integration is stored in $q_{1}$.
V-E Unified Memory
The unified memory approach uses unified-shared-memory (USM) mode on the MI300A (El Capitan), AMD InfinityFabric for CPU–GPU connections on Frontier, and the NVLink C2C connection on Grace Hopper (Alps), Figure˜4 depicts this strategy and uses the full capacity of the compute node, expanding beyond the GPU memory capacity without a meaningful performance hit. This allows us to store the intermediate Runge-Kutta stage on the host, reducing GPU memory use by up to a factor of $12/17$.
V-E1 AMD approach for El Capitan
On the AMD MI300A, the CPU and GPU share one physical HBM pool. We compile MFC with OpenMP offloading directives using AMD’s next-generation Flang compiler and OpenMP’s USM mode. All variables have one copy in memory and are accessible from the CPUs and accelerators with no data movement. CCE’s OpenACC implementation is also used for speed comparisons.
V-E2 AMD approach for Frontier
Our approach to Frontier involves using OpenACC or OpenMP, as well as HPE’s CCE or the AMD Flang next-generation compiler, for performance comparison between compilers and unified memory implementations.
With OpenMP, the mapping of the Fortran pointers associated with the device-resident memory within the MFC data structures is omitted. No redundant memory copy is created due to MP’s treatment of the derived type mapping with Fortran pointer components. MFC can also run on Frontier in USM mode, but was found to be less performant on Frontier than a UVM strategy discussed in section˜V-E3 below, and is not used in this work. In sum, we decrease memory use on the host and device without impacting performance.
V-E3 NVIDIA approach for Alps
The time step updates in the Runge–Kutta scheme are rearranged so that only the current sub-step is passed to the right-hand side routine. The buffer holding the previous state is used to update the current Runge–Kutta state. With this rearrangement, we only store one sub-step and the right-hand side buffer in GPU memory. Thus, the intermediary sub-step is always in CPU memory, freeing GPU memory and increasing simulation size without sacrificing performance. The time step updates access Runge–Kutta sub-steps from their physical locations via zero-copy, accessing data from GPU and CPU memory simultaneously through the C2C NVLink. We include flexibility in hosting the IGR temporary variables in either GPU or CPU memory, further reducing the memory footprint from $12/17$ to $10/17$ and scaling to larger problems with little performance impact.
We communicate via CUDA-aware MPI and GPUDirect, avoiding staging in host memory. We allocate the send and receive buffers on the GPU using OpenACC capture to guide MPI in selecting the GPU path. This strategy enables separate memory for the send/receive halo buffers.
(a) FP16/FP32
(b) FP32
(c) FP64
(d) FP64 (Baseline)
Figure 5: Visualization of a three-engine configuration and its plumes using (a) FP16, (b) FP32, (c) FP64 storage, and (d) the baseline numerics. The contours show where the velocity is supersonic relative to the free stream flow, and the initial state is seeded with smooth random noise in all cases. FP32 and FP64 yield visually indistinguishable results. Visual differences in the FP16 case are solely due to the earlier onset of physical flow instabilities, but are still faithfully representative of the flow features. The grid-dependent nature of the baseline shock-capturing approach results in spurious grid-alignment artifacts.
V-F Mixed Precision
This strategy is numerically viable for various compressible flow simulations, as discussed in section˜IV-C. This strategy further doubles the maximum simulation size over state-of-the-art shock capturing methods we compare against, which are numerically unstable at even single precision.
1 $(\rho,\rho\mathbf{u},E,\alpha)\leftarrow\text{Conservative variables}$

### v4

@@
Abstract.
We present an optimized implementation of the recently proposed information geometric regularization (IGR) for unprecedented scale simulation of compressible fluid flows applied to multi-engine spacecraft boosters. We improve upon state-of-the-art computational fluid dynamics (CFD) techniques in terms of computational cost, memory footprint, and energy-to-solution metrics. Unified memory on coupled CPU–GPU or APU platforms increases problem size with negligible overhead. Mixed half/single-precision storage and computation are used on well-conditioned numerics. We simulate flow at 200 trillion grid points and 1 quadrillion degrees of freedom, exceeding the current record by a factor of 20. A factor of 4 wall-time speedup is achieved over optimized baselines. Ideal weak scaling is observed on OLCF Frontier, LLNL El Capitan, and CSCS Alps using the full systems. Strong scaling is near ideal at extreme conditions, including 80% efficiency on CSCS Alps with an 8 node baseline and stretching to the full system.
@@ Abstract.
copyright: ccjournalyear: 2025conference: The International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: The International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3712285.3771783isbn: 979-8-4007-1466-5/2025/11ccs: Applied computing Physicsccs: Computing methodologies Massively parallel and high-performance simulations
1. Justification for ACM Gordon Bell Prize
Largest compressible computational fluid dynamics simulation, exceeding 1 quadrillion degrees of freedom and a factor of 20 beyond state-of-the-art. Enabled by a large-scale optimized use of inviscid (information geometric) regularization. Time-to-solution and energy-to-solution decrease by up to a factor of 4 and 5.4 in double precision, respectively, with greater decreases in single and mixed precision. Code available at https://github.com/MFlowCode/MFC
2. Performance Attributes
Table 1. Summary of Performance Attributes
@@ 2. Performance Attributes
3. Overview of the Problem
The 21st century witnesses a “new space race” (pekkanen2019governing) driven by private companies replacing government agencies in providing launch services. The resulting drop in launch cost enabled many business models, ranging from satellites to space manufacturing (denis2020new). The new emphasis on cost efficiency motivated innovations in rocket design, such as reusing rocket stages and leveraging economies of scale.
@@ 3. Overview of the Problem
Large arrays of small engines create new design challenges. The exhaust plumes of densely packed engines can interact, propelling hot gas toward the rocket base and heating it. This so-called base heating can cause mission failure (mehta2013numerical), mandating heat shields on the rocket base that increase weight and cost.
Mitigating base heating most cost-effectively, in terms of both dollars and weight, requires understanding the mechanism by which engine exhaust is reflected towards the rocket and identifying which parts are most affected. A key difficulty in experimental approaches to understanding plume recirculation is that it depends on numerous parameters, including varying ambient pressure as the rocket traverses the atmosphere and engine thrust vectoring for steering. A detailed flow field characterization under a broad range of conditions is only feasible with numerical simulations. They even allow probing the impacts of changes to the rocket design. Prior work on the numerical simulation of interacting rocket plumes was limited to small numbers (up to 7) of rocket engines and limited resolution (up to $\approx{}10$ million grid points) (mehta2013numerical; wang2024effect; ren2024numerical). Our work addresses this shortcoming by utilizing the latest flagship exascale systems, leveraging their hardware design and coupling novel algorithms, computational methods, and the optimizations they enable. With this, we can simulate the interaction of rocket engine plumes at unprecedented scales.
@@
(1) $\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})$ $=0,$
(2) $\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u}+p\mathbf{I}-\bm{\tau})$ $=\mathbf{0},$
(3) $\frac{\partial E}{\partial t}+\nabla\cdot\left[(E+p)\mathbf{u}-\mathbf{u}\cdot\bm{\tau}\right]$ $=0,$
@@
(4) $p=(\gamma-1)\rho e,\quad\text{for}\quad e\coloneqq E/\rho-\|\mathbf{u}\|^{2}/2$
@@
(5) $\tau_{ij}=\mu\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right)+\left(\zeta-{2}\mu/{3}\right)\delta_{ij}\frac{\partial u_{i}}{\partial x_{j}}.$
Equations˜1, 2 and 3 describe the conservation of mass, momentum, and energy, and $\bm{\tau}$ is the viscous stress tensor and $p$ the pressure. Tracking the mixture ratios of different gases and fluids, as well as their chemical reactions, is a natural extension of the demonstration. For this work and its implications, we focus on eqs.˜1, 2, 3, 4 and 5.
@@
Figure˜1 shows a visualization for a simulation run on CSCS Alps. The visualization shows the interactions between an array of 33 Mach 10 rocket engines organized in an array inspired by that of the SpaceX Super Heavy, each resolved by 600 grid cells across its outlet. The high resolution allows fine-scale details to be represented despite the high Mach number. The simulation of fig.˜1 uses a rectilinear grid of 3.3T cells and ran for 16 hours on 9.2K GH200s (2300 Alps nodes).
Figure 1. Simulation results showing the interacting plumes from an array of 33 thrusters in a configuration inspired by the SpaceX Super Heavy with 16.5 trillion degrees of freedom. (The thrusters themselves are for visualization purposes. We model them through inflow boundary conditions.) :
4. Current State of the Art
4.1. Shock capturing
@@
Shock waves are the primary concern in high-speed compressible fluid dynamics simulations. They arise in numerous natural and man-made phenomena, including supernovae, air, and spacecraft. The velocity and density fields of compressible high-speed flows sharpen over time, eventually forming macroscopic discontinuities or shock waves. However, on the microscopic scale, the gas viscosity balances the steepening of shock waves, resulting in smooth profiles. In practice, this happens on the scale of the mean free path of gas particles, orders of magnitude smaller than the quantities of interest.
@@
Higher-order numerical methods exploit the target solution’s regularity to approximate it with smooth functions, most frequently polynomials. The mean-free path is many orders of magnitude shorter than a realistic computational grid spacing. Thus, shocks appear as discontinuities on the grid scale, and the direct application of higher-order methods leads to Gibbs–Runge oscillations and subsequent simulation failure. Shock capturing modifies either the equation or its discretization to obtain a well-behaved object on the grid scale. It amounts to coarse-graining the microscopic shock and correctly representing its macroscopic effects without resolution at the grid level. The resulting coarse-scale model should preserve smooth grid-scale oscillations due to turbulence, reactions, or acoustic effects.
@@
Artificial viscosity mitigates Gibbs–Runge oscillations at the cost of excessive dissipation of fine-scale features. To remedy the latter, numerous approaches apply artificial diffusivity adaptively, in the vicinity of the shock (cook2004high; guermond2011entropy; mani2009suitability).
@@
Realizing a viscous regularization is challenging in practice. A lack of viscosity creates spurious oscillations, while excessive viscosity dissipates the solution. Common choices, such as the localized artificial diffusivity (LAD) of (cook2004high), attempt to strike this balance. Viscous methods spread the shock over multiple grid points, but the resulting shock profile is not high-order smooth (fig.˜2 (a,i)). This can still lead to Gibbs–Runge oscillations overcoming the regularization and destabilizing solutions. Increasing the shock width requires increasing the strength of the artificial viscosity, which smears true physical features critical to representing the flow (fig.˜2, (b,i)). In the presence of sufficiently strong shocks, the required artificial viscosity affects the CFL numbers of the explicit time steppers considered state-of-the-art for such hyperbolic problems.
@@
Limiters are an alternative to artificial viscosity. They adaptively lower the order of the numerical method at shocks (van1979towards; liu1994weighted). They are more robust but also risk dissipating fine-scale features. Riemann solvers aim to mitigate this problem but add computational cost (toro2019hllc; roe1981approximate; harten1983upstream).
4.2. GPU memory
@@
The evolution of GPU memory over the past decades, in terms of capacity and bandwidth, has been highly beneficial for CFD applications, which exhibit low arithmetic intensity. Thus, performance is limited by memory bandwidth. The spatio-temporal scale separation in fluid flows requires a high resolution, necessitating large amounts of memory to be accessed at high speeds. Both NVIDIA and AMD have blurred the lines between GPU and CPU memory by introducing coherent CPU-to-GPU interfaces: $900\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ bidirectional bandwidth for the NVIDIA Grace Hopper and InfinityFabric at $8\times 72\,$\mathrm{GB}\text{/}\mathrm{s}$\$ for the AMD Trento+4 MI250X configuration of Frontier. These fast interconnects, along with technologies such as unified virtual memory (UVM), allow CPU memory for GPU computations and even (in the case of Grace Hopper) allow the GPU to saturate the host memory bandwidth. Furthermore, AMD introduced the MI300A in LLNL’s El Capitan and Tuolumne, with a single physical HBM pool accessed by both CPU and GPU devices. This strategy eliminates concerns about device and host pointers or memory storage locations.
4.3. Floating point computation
@@
The advent of artificial intelligence has fueled the development of algorithms with reduced and mixed precision, although their adoption in HPC applications remains limited. Most traditional scientific applications rely on FP64 computation and storage. However, recent advances in software and numerical methods demonstrate the viability of sub-FP64 precision for well-conditioned numerical methods and a range of traditional fluid flow problems (karp2024sensitivity). Still, FP64 is the de facto standard for compressible, shock-laden flow simulation and traditional shock-capturing techniques. WENO-based reconstruction methods and approximate Riemann solves, considered state-of-the-art and used as the baseline in this work, involve poorly conditioned operations and thus are not well suited to reduced precision (Brogi2024OpenFOAMPrecision). We elide such issues herein by avoiding numerical shock capturing entirely via IGR.
4.4. Large CFD simulations
@@
Previous Gordon Bell winner (rossinelli201311) performed a 10T grid point CFD simulation on IBM BlueGene/Q. The system’s substantial CPU memory enabled this simulation. However, the wall time required to process such a large simulation makes achievable physical time scales short. More recently, (sathyanarayana2025high) solved a compressible CFD problem of 10T grid points when extrapolating to 100% of Frontier.
@@
The current largest accelerator-based flow simulations are for incompressible flows, which have fewer degrees of freedom. The largest simulation represented turbulence on a 30T point grid using OLCF Frontier (yeung2025gpu). Large time step costs arise from all-to-all communications, which dominate the time-to-solution.
5. Innovations Realized
5.1. The need for scale
@@
Refer to caption :
Figure 2. Inviscid regularization: Localized artificial diffusion (LAD) spreads shocks over a user-defined width (a,i). The resulting curve is not high-order smooth. This can cause methods with a high order of accuracy to develop oscillations and, ultimately, fail. Increasing the width for coarser discretizations yields unphysical and significant dissipation of oscillatory solution profiles (b,i). Information geometric regularization replaces shocks with smooth profiles (a,ii) at the grid scale and preserves oscillatory features (b,ii).
5.2. Shock treatment with information geometric regularization
We avoid the limitations of standard shock-capturing approaches via the first inviscid regularization of the equations, called information geometric regularization (IGR) as recently proposed by Cao and Schäfer (cao2023information). IGR was first derived in the pressureless (infinite Mach number) case, where shocks amount to the loss of injectivity of the flow map $x\mapsto\phi_{t}(x)$ that maps gas particles from their initial position to their position at time $t$. IGR modifies the geometry according to which the flow map evolves, such that particle trajectories $t\mapsto\phi_{t}(x_{0})$ do not cross and instead asymptotically approach each other (see fig.˜3). This preserves the long-time post-shock behavior, prevents the formation of grid-level singularities, and recovers the nominal vanishing viscosity solutions in the limit. This result has been proven rigorously in the unidimensional, pressureless case (cao2024information).
@@ 5.2. Shock treatment with information geometric regularization
(6) $\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})$ $=0,$
(7) $\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u}+(p+\Sigma)\mathbf{I}-\bm{\tau})$ $=\mathbf{0},$
(8) $\frac{\partial E}{\partial t}+\nabla\cdot\left[(E+p+\Sigma)\mathbf{u}-\mathbf{u}\cdot\bm{\tau}\right]$ $=0,$
(9) $\alpha\left[\operatorname{tr}{(\nabla\mathbf{u})^{2}}+\operatorname{tr}^{2}{(\nabla\mathbf{u})}\right]$ $=\frac{\Sigma}{\rho}-\alpha\nabla\!\cdot\!\left(\frac{\nabla\Sigma}{\rho}\right)\!.$
@@
Refer to caption :
Figure 3. Information geometric regularization modifies shocks by changing the geometry according to which the flow map $\phi_{t}$ evolves in time. In the modified geometry, the trajectories of two tracer particles $t\mapsto\phi_{t}(x_{1}),\phi_{t}(x_{2})$ converge in $t$ rather than cross. The regularization strength $\alpha$ determines the rate of convergence. The vanishing viscosity solution is recovered in the $\alpha\rightarrow 0$ limit. Figure adapted from Cao and Schäfer (cao2023information) with author permission.
@@
IGR allows one to bypass shock capturing, instead using a third- or fifth-order accurate finite volume method directly. Lax–Friedrichs numerical fluxes treat the hyperbolic part of the equation. Due to the high Reynolds number of the problem under consideration, we find that a second-order accurate approximation of the derivatives of $\mathbf{u}$ suffices to compute the viscous stress tensor. We reuse these derivatives to compute the left side of eq.˜9 and discretize the elliptic operator on the right using a standard 7-point stencil.
@@
For each computation of the hyperbolic flux, we solve eq.˜9 using up to 5 sweeps of Jacobi or Gauss–Siedel iteration, with the previously computed $\Sigma$ as an initial guess. We use a third-order accurate Runge–Kutta time stepper (gottlieb1998total). For a single species (advected fluid) case, the total number of floating point numbers stored by our scheme is $17N+o(N)$, where $o(N)$ is the number of grid points. This includes $5$ state variable arrays (density, energy, and three momenta), which are the degrees of freedom of the solution per grid cell. We also hold $5$ arrays for the Runge–Kutta sub-step, $5$ arrays for the right-hand side of the discretized PDE system, one array for $\Sigma$, and one for the right-hand side of eq.˜9. An additional copy of $\Sigma$ is required if Jacobi sweeps are used for the iterative solve.
5.3. The algorithm
[…]